OpenStudy (anonymous):

whats the percent of 50 is 12.5

OpenStudy (anonymous):

25%

OpenStudy (anonymous):

50*0.25=12.5

OpenStudy (anonymous):

50* a%=12.5 a%=12.5/50

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

remember original * % = new number